OpenStudy (anonymous):

i need help plz plz , A movie is labeled as being 90 minutes. How many times could you watch the movie in one day? A)9 B)10 C)15 D)16

OpenStudy (anonymous):

plz dont understand

OpenStudy (raden):

90 minutes = 1.5 hours 1day = 24 hours so, it can be 24/1.5 = ...

OpenStudy (anonymous):

ok so its 15

OpenStudy (raden):

24/1.5 = 24/(3/2) = 24 * 2/3 = 48/3 = ... ?

OpenStudy (anonymous):

im not so good in math

OpenStudy (anonymous):

I think its 16 because there are 24 hours in one day and there or 60 minutes in each hour. So you would do 24 times 60 which equals 1440 and divide that by 90 and that equals 16!!

OpenStudy (anonymous):

ok thanks guys for you help

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

i have this one more q A truck transports 7 cars to an auto dealership. Each car has a gasoline tank capacity of 64 liters. What is the total gasoline tank capacity of all of the cars in kiloliters?

OpenStudy (anonymous):

is it multiple choice?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

what are they?

OpenStudy (anonymous):

A)0.448 b)4.48 c)44.8 d)448

OpenStudy (anonymous):

these are the multiple q

OpenStudy (anonymous):

it is A)0.448

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

no problem :)